### v2

# AuthPrint: Fingerprinting Generative Models Against Malicious Model Providers

Abstract
Generative models are increasingly adopted in high-stakes domains, yet current deployments offer no mechanisms to verify whether a given output truly originates from the certified model. We address this gap by extending model fingerprinting techniques beyond the traditional collaborative setting to one where the model provider itself may act adversarially, replacing the certified model with a cheaper or lower-quality substitute. To our knowledge, this is the first work to study fingerprinting for provenance attribution under such a threat model. Our approach introduces a trusted verifier that, during a certification phase, extracts hidden fingerprints from the authentic model’s output space and trains a detector to recognize them. During verification, this detector can determine whether new outputs are consistent with the certified model, without requiring specialized hardware or model modifications. In extensive experiments, our methods achieve near-zero FPR@95%TPR on both GANs and diffusion models, and remain effective even against subtle architectural or training changes. Furthermore, the approach is robust to adaptive adversaries that actively manipulate outputs in an attempt to evade detection.
I Introduction
Recent advances in generative AI have led to the widespread deployment of generative models across various domains, with providers of generative AI services increasingly monetizing their models by offering subscription-based access. However, this rapid adoption has raised serious concerns about the risks posed by these models, particularly in safety-critical domains, such as healthcare and defense, where erroneous model outputs can have disastrous consequences [1].
In response, policymakers are introducing legal frameworks to regulate the use of AI and, in particular, the deployment of generative models. For instance, the European Union’s AI Act mandates independent, periodic audits for “high-risk” AI systems deployed in domains such as healthcare, education, employment, and critical infrastructure [2]. This requirement to pass or be certified by an audit raises a critical question: How can users verify that a given output indeed originated from the audited model? Even if a model passes an audit, there is currently no guarantee that providers will adhere to the audited model and not substitute it for a cheaper but potentially flawed version.
Without verification mechanisms, audits may fail to enforce compliance. In particular, it becomes difficult to detect whether the model deployed after an audit matches the one that had been audited. Similar challenges have been observed in other sectors, where companies have deceived auditors by manipulating a product’s performance only to pass an audit, while deploying a product with detrimental effects on climate and public health [3]. In the context of AI, service providers have similar intentions to reduce costs, possibly disregarding harmful societal outcomes [4].
This issue is exacerbated in current AI/ML systems, where access to the model is offered through opaque APIs that provide no verifiable information about which models are accessed. Users are left to blindly trust the model provider: for any single query, the model provider can always plausibly deny having used a harmful or defective model, as users have no access to evidence that could disprove these claims. This lack of accountability hinders oversight and diffuses responsibility across the ML supply chain [5, 6]. Therefore, technical means to track models through their lifecycle—from certification to deployment—is critical, not only for law enforcement but also for maintaining trust in AI applications.
Several cryptographic methods have been proposed in the literature to verify that outputs originate from a specific model. While early cryptographic verification protocols were highly interactive and thus not applicable on single queries [7], recent breakthroughs in the field of zero-knowledge have enabled proofs that can verify that a previously committed model is evaluated on a given query [8]. Recent zk-ML systems compile small networks into arithmetic circuits to enable zk-SNARKs [9, 10, 11, 12, 13], but the overheads of these systems—in terms of the added latency of proof verifications— still limit their practical feasibility, particularly in domains where fast responses are critical.
Another line of work binds models to trusted execution environments (TEEs), with outputs verified via remote attestation [14]. Yet GPU-based TEEs remain immature, and hybrid schemes that delegate computations to untrusted GPUs require weight quantization, adding latency and reducing accuracy. Overall, both cryptographic and TEE-based solutions face scalability, efficiency, and hardware barriers, making them impractical in many deployments of large generative models.
In this paper, we propose AuthPrint (Authentication via Fingerprinting), a novel framework for detecting the source model of generated images. AuthPrint relies on a trusted third party (e.g., an auditor) to train a model that learns to reconstruct secret fingerprints from images generated by the original audited model. During verification, the reconstructed fingerprint is compared against the true fingerprint to determine whether the image was produced by this authentic model.
Crucially, our empirical evaluation shows that AuthPrint satisfies two essential properties: (i) it reliably detects deviations in output distributions from the original audited model, and (ii) it remains robust against attacks aimed at forging images that pass verification. We show that AuthPrint achieves consistently low detection error on both StyleGAN2 and Stable Diffusion models, even under model substitutions with minor training modifications, model compression, and adaptive adversarial attacks based on evasion and fingerprint recovery. These results position AuthPrint as an effective method for authenticating image generation models across diverse architectures and adversarial settings.
To sum up, our main contributions are:
  • We introduce AuthPrint, a black-box covert fingerprinting framework for attributing image outputs to a specific generative model, even in the presence of a malicious model provider.
  • We demonstrate its effectiveness across diverse architectures and datasets, including StyleGAN2 (trained on FFHQ and LSUN-Cat) and multiple versions of Stable Diffusion.
  • We show that AuthPrint is robust to adaptive adversaries with full access to the original audited model, reducing forgery success rates from 100% (in baseline methods) to zero.
  • We provide a theoretical analysis of fingerprint recovery attacks, establishing their infeasibility under realistic adversarial constraints.
II Background
In this section, we define the problem and introduce the necessary background on model fingerprinting techniques.
II-A Problem Statement
We consider a setting involving two entities: the model provider and the verifier. The model provider maintains a generative image model and monetizes it by offering remote access (e.g., via a web API). We assume the model has been trained and can be represented as a function mapping structured inputs to high-dimensional image outputs. More formally, we define an image generation model as a function $G_{\psi}:\Xi\times\mathcal{C}\to\mathbb{R}^{d}$, where $\psi$ denotes its parameters, $\Xi$ is the space of latent variables, $\mathcal{C}$ is the space of conditioning inputs, such as class labels or text prompts, and $\mathbb{R}^{d}$ is the output space corresponding to images of dimension $d=\text{channels}\times\text{height}\times\text{width}$. The parameters $\psi$ include everything required to fully specify the model for inference. For example, if $G_{\psi}$ is implemented as a neural network, $\psi$ would include its weights and architecture.
Before deployment, the model provider must grant the verifier query access. We refer to this pre-deployment stage as the certification phase. During this phase, the verifier may issue a number of queries via the API, and the provider responds by evaluating $G_{\psi}$ on the verifier’s inputs and returning the resulting output images. We refer to the period after the model has been deployed as the verification phase. The central question we investigate is: Given any output $x\in\mathbb{R}^{d}$ at verification time, can the verifier determine whether it was produced by the model that was certified?
We envision real-world scenarios where this problem is relevant. Our running example is an AI audit, where the verifier is an auditor certifying the model’s compliance with AI regulations, such as the need to meet appropriate fairness and privacy requirements for the specific application. Depending on how the regulations are implemented, the provider may grant the auditor query access, or provide full access to the model under legal non-disclosure agreements. Next, we detail the goals and capabilities of the model provider and verifier.
II-B Threat Model
We consider a model provider aiming to maximize profit from its image generation service. The main threat we address is post-certification model substitution: the provider may deploy, behind an opaque API, a different generator of inferior quality but cheaper to evaluate. Such a bogus model can introduce flaws that are imperceptible in single images yet material for downstream use (e.g., degraded medical image fidelity). In this work we focus on distributional shifts that manifest as quality deterioration, while noting that other shifts (e.g., privacy leakage or demographic skews) are also plausible.
Certification-phase authenticity assumption. Our setting targets deployments subject to oversight (e.g., regulatory or contractual audits) in which the provider grants a verifier controlled query access before deployment. During this certification phase, the verifier interacts with the provider’s model instance to build a detector $D$. We assume the instance exposed at certification is the authentic model intended for deployment. This mirrors standard audit practice: access occurs under legal or commercial obligations that penalize substitution at certification time. The assumption is operationally realistic and can be reinforced with lightweight safeguards during certification, such as session binding or simple challenge–response checks, though these are not required at verification. Our contribution then addresses the post-certification risk: once the model is deployed behind an opaque API, $D$ must enable users to verify that outputs still come from the certified distribution.
Verifier’s goal and capabilities. The verifier that we consider has only black-box access during certification but can issue many queries to the authentic model to train a detector $D$ that, given an image $x$, decides whether $x$ was sampled from the certified model’s output distribution. The verifier does not alter the provider’s model and does not rely on specialized hardware. After deployment, the verifier exposes $D$ as a separate black-box verification API that returns only a binary output about the authenticity of the queried model.
Adversary’s goal. After certification, the provider may replace the authentic generator with a cheaper substitute or a modified variant and still have outputs accepted by $D$. Concretely, the adversary aims to produce $\hat{x}\in\mathbb{R}^{d}$ whose distribution deviates from that of the certified model while $D$ incorrectly accepts it as authentic (a forgery attack).
Adversary’s knowledge and capabilities. During certification, the provider must answer queries using a single fixed model instance; attempting a swap during the certification phase can be deterred by oversight and additional safeguards. Post-certification, the provider has full white-box access to the authentic model. However, the provider has no access to the internals of the verifier’s private detector and can only see the binary accept/reject outputs of their queries—which are likely to be rate-limited by the verifier.
We assume the adversary is not constrained in the architecture or training choices they can make to develop the substitute model, except that the cost of evaluating the substitute model should not exceed that of the authentic model, as the motivation for substitution is cost saving. This rules out attacks that are systematically more expensive than sampling from the certified generator.
In summary, we assume authentic access only during a controlled, pre-deployment certification window—an assumption aligned with real audit workflows—and we design $D$ to detect subsequent model substitution when users can see only the provider’s opaque API outputs.
Our approach to this problem is based on model fingerprinting techniques, which we introduce next.
II-C Model Fingerprinting Techniques
Model fingerprinting refers to methods that identify machine learning models by exploiting distinctive statistical patterns in their outputs. These methods are typically motivated by intellectual property protection: when a model is stolen, leaked, or re-deployed without authorization, its fingerprint can be used to trace the source model.
Fingerprinting techniques are commonly divided into two categories: active and passive. Active fingerprinting modifies the model itself—for example, by embedding secret perturbations in the weights—so that its outputs carry an identifiable signal. While effective in cooperative or white-box settings, active methods are fragile when the provider may behave adversarially: access to both the fingerprinted and unmodified models allows side-by-side comparisons to expose or strip the fingerprint.
By contrast, passive fingerprinting does not alter the model, but instead relies on statistical regularities that naturally emerge in its outputs. This makes passive approaches better suited for our setting, where the verifier only has black-box access to the model during certification.
Passive fingerprinting for model attribution. A growing body of work has explored passive model fingerprinting techniques for attributing generated images to their source model. Yu et al. demonstrate that noise residuals of generated images contain generator-specific traces, and propose deep architectures that exploit these patterns to identify which of several GANs produced a given image [15]. Dzanic et al. show that generative models leave distinctive patterns in the Fourier spectrum—particularly in the high frequency band—that can serve as reliable fingerprints to distinguish different models [16]. Girish et al. extend attribution to scenarios where new generators may emerge post-training, by combining out-of-distribution detection and clustering to discover and assign samples from previously unseen models [17]. More recently, Song et al. take a geometric approach in which generated samples are projected onto the manifold of real images, with the resulting offsets serving as model-specific fingerprints [18]. Follow-up work generalizes this approach to Riemannian geometry, yielding more faithful representations of data structure and improved attribution across diverse datasets and architectures [19].
These studies typically report model attribution accuracies in the range of 70–99% with a pool of between 5 to 20 candidate models. Although they are evaluated in small worlds of models, the techniques show promise for reliable model identification, aligning with the objectives of our work. However, these techniques are not suitable to our setting: we consider an adversarial setting where an untrusted provider may substitute the certified model after deployment to reduce costs or evade oversight. In this case, the adversary has full access to both the audited and substitute models, and the fingerprinting techniques. Therefore, if these fingerprinting methods were used, the adversary could easily replicate or mimic the fingerprints, thereby bypassing detection.
To our knowledge, no existing passive fingerprinting approaches have been designed to address such an adversarial setting, nor have they been evaluated under adaptive forgery attacks. Furthermore, prior evaluations often compare across heterogeneous model families (e.g., different GAN architectures or distinct generator types), whereas our setting requires the techniques to be robust against far subtler modifications, such as variants trained with slightly different configurations or models subjected to compression techniques. These gaps motivate the approach we introduce in the next section.
III AuthPrint: Authenticity Auditing via Covert Model Fingerprinting
We introduce AuthPrint (Figure 1), a covert, passive fingerprinting framework for verifying the outputs of image generation models when providers of these models may behave adversarially. AuthPrint is run by the verifier to build a fingerprint detector during the certification phase and later use it to determine the authenticity of model outputs during the verification phase.
AuthPrint could be offered as a service to users of the model provider’s API, in which case the verifier operates as a trusted third party. In many real-world scenarios, there are trusted entities who could play the role of the verifier, such as regulatory agencies or public interest auditors.
Figure 1: Overview of the AuthPrint pipeline. In the certification phase, a verifier with black-box access to the claimed model selects secret pixel locations (image channels are simplified for illustration) as a fingerprint and trains a private reconstructor to predict them from generated images. After certification, the provider deploys the model via its black-box API, while the verifier hosts a separate verification API with the reconstructor and secret indices. In the verification phase, a user submits an image to this API, which applies the reconstructor, compares predicted and actual values at the secret locations, computes a detection error, and returns the authenticity verdict.
Certification Phase. As shown in Algorithm 1, the verifier is granted black-box access to a generative model $G_{\psi}$ and its input distribution $P_{\mathrm{in}}$. This input distribution defines a joint prior over latent variables and optional conditioning inputs, i.e., $(\xi,c)\sim P_{\mathrm{in}}$, where $\xi\in\mathbb{R}^{k}$ is typically sampled from a normal distribution, and $c$ represents conditioning information, such as prompts. The verifier selects a secret sequence of $l$ pixel indices uniformly at random as a vector, $s\in[d]^{l}$, where $d$ is the image dimension. The vector $s$ constitutes the model’s fingerprint which, importantly, is output-dependent.
Algorithm 1 Certification Phase
[…]
Parameter: Fingerprint length $l$, batch size $B$, learning rate $\eta$, training steps $T$
Output: Detector $D$
1: Sample secret index vector: $s\overset{\mathdollar}{\leftarrow}\mathcal{U}([d]^{l})$
2: Initialize reconstructor parameters $\phi$
3:for $t=1$ to $T$ do
4:  Sample batch $\{(\xi_{i},c_{i})\}_{i=1}^{B}\overset{\mathdollar}{\leftarrow}P_{\mathrm{in}}$
5:  Generate images: $x_{i}\leftarrow G_{\psi}(\xi_{i},c_{i})$
6:  Reconstruct fingerprint: $r_{i}\leftarrow R_{\phi}(x_{i})$
7:  Extract fingerprint: $f_{i}\leftarrow(x_{i})_{s}$
8:  Compute loss: $\mathcal{L}\leftarrow\frac{1}{B}\sum_{i=1}^{B}\frac{1}{l}\|r_{i}-f_{i}\|_{2}^{2}$
9:  Update reconstructor: $\phi\leftarrow\phi-\eta\,\nabla_{\phi}\mathcal{L}$
10:end for
11:return detector $D=(R_{\phi},s)$
The goal of the certification phase is to train a fingerprint reconstructor model $R_{\phi}:\mathbb{R}^{d}\to\mathbb{R}^{l}$—typically a deep neural network—that can predict the fingerprint values (i.e., the pixel values at the secret indices $s$) directly from a generated image. To train the reconstructor, the verifier repeatedly samples batches $(\xi_{i},c_{i})$ from $P_{\mathrm{in}}$ and generates corresponding images $x_{i}=G_{\psi}(\xi_{i},c_{i})$. For each image, the reconstructor computes a prediction $r_{i}=R_{\phi}(x_{i})$, and the true fingerprint $f_{i}=(x_{i})_{s}$ is extracted according to $s$. The fingerprint reconstructor is trained to minimize the mean squared error (MSE) between prediction and ground truth fingerprint values over the batch, i.e., $\mathcal{L}=\frac{1}{B}\sum_{i=1}^{B}\frac{1}{l}\|r_{i}-f_{i}\|_{2}^{2}$, where $B$ is the batch size. After training, the verifier outputs a fingerprint detector $D=(R_{\phi},s)$, consisting of the reconstructor $R_{\phi}$ and the secret indices $s$ used for training $R_{\phi}$.
Verification Phase. As shown in Algorithm 2, once trained, the detector $D=(R_{\phi},s)$ is deployed as a black-box verification API controlled by the verifier. At inference, a user submits a test image $x^{\prime}$; the reconstructor computes $r^{\prime}=R_{\phi}(x^{\prime})$, and the verifier also extracts the fingerprint $f^{\prime}=(x^{\prime})_{s}$. The detection error $e=\frac{1}{l}\|r^{\prime}-f^{\prime}\|_{2}^{2}$ is compared to a pre-agreed threshold $\tau$ (set by the provider and verifier to yield low FNR). If $e<\tau$, the image is deemed authentic, meaning it originates from $G_{\psi}$; otherwise, not authentic.
Algorithm 2 Verification Phase
Input: Test image $x^{\prime}$, detector $D=(R_{\phi},s)$
Parameter: Threshold $\tau$
Output: Binary decision
1: Reconstruct fingerprint: $r^{\prime}\leftarrow R_{\phi}(x^{\prime})$
2: Extract fingerprint: $f^{\prime}\leftarrow(x^{\prime})_{s}$
3: Compute detection error: $e\leftarrow\frac{1}{l}\|r^{\prime}-f^{\prime}\|_{2}^{2}$
4:if $e<\tau$ then
5:  return Authentic (i.e. $x^{\prime}$ originates from $G_{\psi}$)
6:else
7:  return Not authentic
8:end if
The core intuition behind AuthPrint rests on two factors. First, modern generative models exhibit fine-grained pixel-level dependencies [20, 21], which enable a reconstructor (a deep neural network) to accurately recover a secret subset of pixels from the full image. Second, we exploit the out-of-distribution (OOD) sensitivity of deep neural networks: the reconstructor performs reliably only when images are sampled from the certified model’s distribution but fails under substitution, as even subtle shifts in conditional dependencies introduce prediction errors. This approach to detecting distribution shifts provides a principled mechanism for determining whether outputs originate from the claimed model.
To instantiate the fingerprint, we sample the index vector $s$ uniformly at random. Random selection ensures that the fingerprint probes pixel dependencies without bias toward specific regions or channels and avoids predictable structures that an adversary might exploit. It also maximizes adversarial uncertainty: if $s$ were fixed or exhibited any discernible pattern, a provider could evade detection by training a substitute reconstructor for gradient-based attacks or by strategically adjusting the chosen pixels. By keeping $s$ both random and secret, any successful forgery must implicitly preserve the full joint distribution of pixels, since the adversary cannot anticipate which subset will be verified.
The security of AuthPrint therefore rests on the secrecy of the detector, namely the reconstructor parameters $\phi$ and the index vector $s$. Although the model provider knows the verification protocol and may interact with the detector, they only observe binary outcomes. Without access to logits, gradients, or the indices themselves, they cannot train surrogates or mount gradient-based evasion attacks. As we show in later sections, recovering $s$ reduces to a threshold group-testing problem to identify the fingerprint set, followed by factorial search over all permutations to recover its order—an infeasible task at practical fingerprint lengths. This information asymmetry, combined with the unpredictability of random index selection, underpins AuthPrint’s robustness against forgery attacks with either evasion or reverse-engineering attempts.
IV Experimental Setup
We evaluate AuthPrint on two representative families of image generators: GAN-based and diffusion-based models, instantiated by StyleGAN2 and Stable Diffusion (SD), respectively. Our evaluation spans a range of adversarial behaviors, including model substitution, model modification, and adaptive attack scenarios. This section provides details on the models, datasets, training configurations, and attack setups.
Generative Models.StyleGAN2: We use models from the official StyleGAN2-ADA repository by NVLabs [^1] https://github.com/NVlabs/stylegan2-ada-pytorch, trained on FFHQ and LSUN-Cat, each producing $256\times 256$ images. For each dataset, we designate one target model and treat variants trained under different configurations as negative models:
  • FFHQ:
    • Target model: trained on 70k images with ADA (Adaptive Discriminator Augmentation) [22].
    • Negative models: trained on
        1k images with ADA
        30k images with ADA
        70k images with BCR (Balanced Consistency Regularization) [23]
        70k images with no augmentation
  • LSUN-Cat:
    • Target model: trained on 100k images with ADA.
    • Negative models: trained on
        1k images with ADA
        30k images with ADA
        100k images with BCR
        100k images with no augmentation
ADA is an adaptive augmentation strategy applied during discriminator training [22], while BCR is a regularization method that enforces consistency across augmented samples [23]. In both datasets, negative models differ in training set size or augmentation strategy, resulting in subtle but detectable distribution shifts.
Stable Diffusion: We use SD 2.1 as the target model and earlier versions (SD 1.5–1.1) as negatives, obtained from HuggingFace [^2] https://huggingface.co. Unless otherwise noted, these models generate $1024\times 1024$ images.
Fingerprint Reconstructor. The reconstructor is a CNN trained to predict pixel values at a secret subset of locations given the full generated image. The architecture of the CNN follows a conventional design: a convolutional encoder with stride-2 downsampling, $4\times 4$ kernels, batch normalization, and LeakyReLU activations, followed by global average pooling and an MLP head. For StyleGAN2 experiments, we use a 32M-parameter reconstructor, which is sufficient given the relatively low diversity and resolution of StyleGAN2 outputs. In contrast, SD experiments employ a larger 674M-parameter model to better capture the higher complexity and semantic richness of text-to-image generations. To study the effect of scaling, we additionally evaluate reconstructors with 32M, 187M, and 674M parameters on SD.
Training Data. For StyleGAN2, training data are generated by sampling latent vectors $z\sim\mathcal{N}(0,I)$. For SD, we default to a fixed, high-specificity prompt—referred to as the Japanese café prompt—chosen for its repeatability and richness of visual detail. This prompt was generated using OpenAI’s ChatGPT-4o model [24]:
[…]
Detection error. We quantify the performance of AuthPrint in detecting the audited model by measuring detection error as the FPR at 95% TPR, treating samples of the targeted model as instances of the positive class and samples of the other models as instances of the negative class. For simplicity, we report detection error results for a single run throughout the paper, as the variance across random seeds was consistently below 1%.
Generation and Training Setup. SD images are generated with 25 denoising steps, classifier-free guidance scale 7.5, and fp16 precision. Reconstructors are trained with Adam optimizers to minimize MSE between predicted and ground-truth fingerprint values. Learning rates are set to $10^{-3}$ for StyleGAN2 and $10^{-4}$ for SD.
Prompt Specificity Study. To study the role of prompt conditioning in guided models, we evaluate 5 prompt regimes for SD spanning different levels of prompt specificity:
  1. I.
    Diverse prompts from DiffusionDB [25],
  2. II.
    “People”-category prompts from PartiPrompts [26],
  3. III.
    A focused subset of simple “People” prompts (e.g., “a man”, “a woman”, “a child”),
  4. IV.
    A single fixed generic prompt (“a high quality photo”),
  5. V.
    A single fixed specific prompt (Japanese café)
These regimes progressively constrain the output distribution, allowing us to assess how prompt specificity impacts detection.
Forgery Attacks. To evaluate evasion-based forgery, we apply projected gradient descent (PGD) attacks [27] against a detector whose reconstructor is trained on a StyleGAN2 target model (FFHQ-70k-ADA) with fingerprint length 32. Negative classes are derived from four model variants: 1k-ADA, 30k-ADA, 70k-BCR, and 70k-no-aug. The PGD attack is configured with an $\ell_{\infty}$ budget of $\epsilon=0.5$, step size 0.001, momentum 0.9, and a maximum of 500 iterations, with early stopping once a successful forgery is found. Since we assume a threat model in which the detector is hidden from the adversary, we simulate the attack using a CNN classifier (with ResNet-18 structure [28]) trained to discriminate between each negative distribution and the target distribution. The gradient of this surrogate ResNet-18 is then used to guide PGD in perturbing images so as to deceive the true detector. We compare this detector’s robustness to two baseline detectors, described in the corresponding results section.
Software and Hardware. All experiments are implemented in PyTorch and executed on NVIDIA H200 GPUs. Training an AuthPrint pipeline on a StyleGAN2 model typically requires 4-6 GB GPU memory, while for SD it requires 17-23 GB GPU memory.
Reproducibility. Following open science best practices, we will publish a repository with the documentation and code for all our experiments upon acceptance of the paper. The repository will include our implementation of AuthPrint, data processing helper methods, and the datasets used. We will also provide detailed instructions on how to reproduce the results of our experiments.
V Results
This section provides a comprehensive empirical evaluation of AuthPrint, organized around key design choices, scalability, and robustness. We first analyze the role of fingerprint length in detection accuracy across GAN- and diffusion-based generators under model substitution. We then examine the effects of scaling training resources—specifically reconstructor capacity and dataset size—on performance, and evaluate ensemble detection as a means of mitigating false positives. Next, we assess AuthPrint under model modifications such as pruning and quantization, and investigate the influence of prompt specificity in conditional generative models like SD. Finally, we study adversarial robustness, focusing on forgery attacks instantiated via evasion and fingerprint recovery.
Our findings reveal several consistent trends. First, appropriate tuning of fingerprint length is critical: intermediate lengths achieve the best balance between signal strength and reconstructability, yielding consistently low detection error (e.g. 0 for StyleGAN2 and near-0 for SD). Second, detection performance on diffusion models improves markedly with larger reconstructors and increased training samples, underscoring the importance of scalability in narrowing the gap with GANs. Third, ensemble detection and prompt specificity both contribute to stronger performance. Finally, AuthPrint demonstrates resilience to adaptive adversaries under realistic threat models, with security deriving from the secrecy of fingerprint indices and the inherent difficulty of reconstructing ordered fingerprints under a black-box setting. Collectively, these results establish AuthPrint as a practical and reliable method for model authentication across diverse generative architectures and adversarial conditions.
V-A Effect of Fingerprint Lengths
To investigate the effect of fingerprint length on AuthPrint’s detection performance, we first conduct an ablation study by measuring AuthPrint’s detection error across varying fingerprint lengths, while holding all other factors constant. We conduct this ablation study for both families of models: GAN-based (StyleGAN2) and diffusion-based (SD) models (see Section IV for details).
Figure 2 shows the detection error for StyleGAN2 models, evaluated on FFHQ (left) and LSUN-Cat (right). Fingerprint lengths increase in powers of 2 to cover the model’s output space. StyleGAN2 target and negative models share the same architecture but differ in training configurations such as dataset size and data augmentation. Thus, this setup probes AuthPrint’s sensitivity to subtle changes in the training procedure that induce visually imperceptible variations in image quality.
Figure 3 shows the results of the evaluation for SD models, where the target model is SD 2.1, and negatives are earlier versions (SD 1.5 to 1.1), introducing progressively larger distributional shifts due to architectural and dataset differences. The maximum fingerprint size is larger compared to StyleGAN2 due to the larger output size of SD models.
Figure 2: AuthPrint detection performance on model substitution for pre-trained StyleGAN2 models (left: FFHQ, right: LSUN-Cat) across varying fingerprint lengths (number of secret pixels chosen). Target and negative models share the architecture but differ in training data size or augmentation, as detailed in the figure. The $x$-axis is in log-scale.
As seen in the figures, AuthPrint consistently achieves low detection error across both model families, demonstrating robustness in separating target from non-target outputs. Detection performance peaks at intermediate fingerprint lengths: 16–32 for StyleGAN2 and 64 for SD. Very short fingerprints provide too little signal for the reconstructor to reliably recover the fingerprint, while excessively long fingerprints create redundancy that dilutes useful information and reduces accuracy under fixed model capacity. These results suggest that there is a sweet spot in fingerprint length, where the signal is rich enough to capture model-specific dependencies but compact enough to be reconstructed effectively. In practice, this implies that the verifier can tune fingerprint length to find an optimal accuracy.
Figure 3: AuthPrint detection performance on model substitution for SD models across varying fingerprint lengths (number of secret pixels chosen). The target is SD 2.1; negative models are earlier versions (SD 1.5, 1.4, 1.3, 1.2, 1.1) with lower image qualities. The Japanese café prompt is used throughout. The $x$-axis is in log-scale.
[…]
Another factor contributing to the performance gap is the generative process itself. GANs often leave subtle artifacts or statistical fingerprints that AuthPrint may implicitly exploit [29]. Diffusion models, by contrast, progressively suppress such artifacts through iterative denoising during the generation process. Figure 4 illustrates this effect by comparing detection results under different inference schedules—25, 15, and 5 denoising steps—where each “step” corresponds to one iteration of the reverse diffusion process during image generation. The AuthPrint reconstructor is trained on SD 2.1 as the target model and evaluated against five negative models (SD 1.1–1.5). Across all negatives, reducing the number of denoising steps consistently lowers detection error, suggesting that shallower sampling may retain more model-specific traces for SD models that aid detection.
Figure 4: Impact of the number of denoising steps (i.e. iterations of reverse diffusion) on the performance of AuthPrint for SD models. The evaluation uses the Japanese café prompt, with a fingerprint length of 1024, a reconstructor with 674M parameters, and 512k training samples. We compare three inference settings: 25, 15, and 5 steps.
V-B Scaling Up Training Enhances Detection
Motivated by the performance gap between StyleGAN2 and SD models, we examined whether increased training resources improve detection. We focused on two factors: (1) reconstructor model size and (2) number of samples for training the reconstructor, both of which significantly affected detection performance.
Figure 5: Effect of reconstructor size (left) and reconstructor training set size (right) on AuthPrint performance for SD models. Both experiments use the Japanese café prompt and a fingerprint length of 1024. Left: reconstructor size comparison (32M, 187M, 674M) using 512k training samples. Right: training set size comparison (128k, 320k, 512k samples) using a 674M-parameter reconstructor.
As shown in Figure 5, we evaluated reconstructors with 32M, 187M, and 674M parameters on SD models (left subfigure), holding all other variables constant. The smallest (32M) failed to reject negative samples, indicating insufficient capacity for the fingerprinting task. In contrast, performance improved substantially with larger reconstructors, and the 674M model achieved near-perfect detection error on SD 1.1.
We also varied the reconstructor’s training set size (128k, 320k, 512k samples) and observed vanishing detection error with more data (right subfigure). These results show that scaling both reconstructor capacity and training volume significantly improves performance. In our most demanding setting (674M reconstructor, 512k samples), training took $\sim$240 GPU hours on an NVIDIA H200—costly for academic clusters but feasible in industrial or governmental settings. Importantly, training is only needed once per target model.
[…]
V-C Boosting Performance with Ensemble Detection
To reduce FPR, AuthPrint can be naturally extended with an ensemble detection scheme, where multiple reconstructors are independently trained using different fingerprint index vectors. At verification, each image is evaluated by all reconstructor-index pairs, and the final detection error is defined as the worst-case MSE across the ensemble, i.e., $e(x^{\prime})=\max_{i\in[N]}\text{MSE}(R_{\phi_{i}}(x^{\prime}),(x^{\prime})_{s_{i}})$, where each $R_{\phi_{i}}$ is trained on $s_{i}$, and $N$ is the ensemble size. This conservative strategy boosts sensitivity: while a negative sample may yield low error on one reconstructor-index pair, it is unlikely to do so across all pairs, thereby reducing false acceptance rates.
Figure 6: Effect of ensemble detection on AuthPrint performance for SD models using the Japanese café prompt. Each reconstructor has 674M parameters, trained on 512k samples with fingerprint length 64.
We evaluate this scheme on SD models using identical hyperparameters but distinct index sets per reconstructor. As shown in Figure 6, detection error decreases with ensemble size, with diminishing returns suggesting convergence. The ensemble design is flexible and can be extended to heterogeneous configurations with varying fingerprint lengths, architectures, or training setups. We leave exploration of the parameters of ensemble approaches to future work.
V-D Detection of Target Model Modifications
We evaluate AuthPrint’s sensitivity to model compression (weight quantization and magnitude-based pruning) and output image downsampling. Quantization and pruning reflect model modifications that a malicious model provider might conduct to save costs at the expense of performance, while advertising the original, unmodified model. Although downsampling is not strictly a model modification and falls outside our threat model—since it requires a postprocessing step after image generation and thus incurs additional costs to the provider—we include it to study the impact of image quality degradation on detection error. We experiment with varying levels of pruning, quantization, and downsampling.
Figure 7: Effect of model modifications on detection performance for StyleGAN2 FFHQ at fingerprint length 32. We evaluate model compression (quantization and pruning) and output perturbations (downsampling). Blue dots (left y-axis) indicate FPR@95%TPR; red dots (right y-axis) indicate FID. Each plot varies the quantization level, pruning ratio, or downsampling resolution to show trends.
Figure 7 shows the relationship between detection error (FPR) and image quality (FID) under all optimizations. AuthPrint achieves effective trade-offs under pruning and downsampling. For example, moderate pruning (62.5%) increases FID modestly to 14.69 but drives FPR to near-zero. Similarly, downsampling from $\times256$ to $\times224$ sharply reduces FPR despite only slight increases in FID. Quantization, by contrast, appears more robust to detection: high precision loss (INT4) renders high FID (>> 200) and leads to 0 FPR as expected, but moderate quantization (INT6) results in high FPR despite quality degradation (FID $\approx$ 53). These findings suggest that while AuthPrint is highly effective at detecting distribution shifts induced by pruning and downsampling, quantization presents a greater challenge. With appropriate tuning of fingerprint length and other pipeline hyperparameters, or through ensemble strategies, deployment in real-world settings may be made even more robust against model compression techniques. This opens a promising avenue for future investigations.
V-E Prompt Specificity Enhances Detection for SD
Unconditioned generators like StyleGAN2 sample from a fixed latent distribution, yielding a well-defined and bounded output distribution $p(x)$. In contrast, conditional models such as SD are typically trained to solve more general generation tasks, conditioned to a prompt $c$, and thus have a much broader range of output distributions $p(x\mid c)$.
This raises a key question for AuthPrint: How does prompt specificity affect detection performance? We evaluate detection error under five prompt regimes, ordered by increasing specificity (see Section IV for prompt details):
  1. I.
    Diverse prompts from DiffusionDB,
  2. II.
    “People”-category prompts from PartiPrompts,
  3. III.
    A focused subset of simple “People” prompts (e.g., “a man”, “a woman”, “a child”),
  4. IV.
    A single fixed generic prompt (“a high quality photo”),
  5. V.
    A single fixed specific prompt (Japanese café)
Figure 8: Effect of prompt specificity on detection performance for SD. We compare five regimes: I. Diverse prompts from DiffusionDB; II. Category-specific “People” prompts from PartiPrompts; III. A focused subset of II (see Section IV); IV. A single generic prompt (“a high quality photo”); V. A single specific prompt (Japanese café prompt).
As shown in Figure 8, detection improves consistently with greater prompt specificity. For diverse prompts, FPR exceeds 80%, indicating poor separation between models. Restricting prompts to the “People” category reduces FPR to below 65%, and the focused subset lowers it to around 10%. The best performance (FPR << 5%) occurs under a fixed, specific prompt. A fixed generic prompt performs similarly in some cases, but with slightly higher error overall.
These results follow a pattern: diverse prompts lead to a broad, overlapping output space, making it harder for the detector to distinguish models. Specific prompts constrain the output distribution, making model differences more detectable. Importantly, a high performance only for specific prompts does not limit practical applicability. Many real-world applications—e.g., profile portraits, product photos, architectural renderings, or avatars—rely on consistent, repeatable prompts, where our approach can be effective.
V-F Evaluation of Adaptive Adversarial Attacks
V-F1 Evasion Attacks
A common strategy for bypassing detectors is evasion, typically implemented via gradient-guided perturbations such as Projected Gradient Descent (PGD) [27]. In our context, an evasion attack corresponds to forgery: an adversary perturbs an image sampled from a negative (non-target) generator so that the detector classifies it as coming from the target model.
We evaluate this threat in a model-substitution scenario. The AuthPrint detector’s reconstructor is trained on a StyleGAN2 target model (FFHQ-70k-ADA) with fingerprint length 32. Negative samples are drawn from four substitute StyleGAN2 variants (1k-ADA, 30k-ADA, 70k-BCR, and 70k-no-aug), which—when the reconstructor is trained on the target—each achieve 0% FPR at 95% TPR (see Figure 2, left). In this threat model we assume the adversary has full access to the target and negative generators but only black-box interaction with the detector itself.
We compare the robustness of three detectors:
  1. 1.
    Baseline: A ResNet-18 classifier trained directly to distinguish target outputs from negatives (no fingerprinting).
  2. 2.
    Yu et al. (2019): A state-of-the-art, public, output-independent model-fingerprinting method [15].
  3. 3.
    AuthPrint: Our method, which uses a reconstructor trained with secret, output-dependent fingerprints.
Because the detector is assumed inaccessible, the adversary performs a surrogate-based, black-box attack: they train a surrogate detector and use its gradients to guide PGD on inputs sampled from the negative generators. Concretely, for each of the three detector designs we train a surrogate ResNet-18 classifier to distinguish the corresponding negative distribution from the target distribution, then run PGD using the surrogate’s gradients to craft perturbations. PGD hyperparameters are given in Section IV.
For each detector we report three quantities: (i) the attack success rate (fraction of negative images that the attack converts to be accepted as target), (ii) the average number of PGD steps to success, and (iii) the LPIPS perceptual distance between original and forged images. These metrics jointly capture how effective the forgery is and how perceptible the required perturbation is.
TABLE I: PGD-based evasion attack results under model substitution. Three detectors are evaluated: (i) Baseline: a baseline detector (ResNet-18), (ii) Yu-2019: a model fingerprinting-based detector proposed by Yu et al. [15], (iii) AuthPrint: our method. Each group reports: attack success rate (ASR, %), average PGD steps, and LPIPS distance for successfully attacked images. “–” indicates no successful attack within 500 steps. The positive model is FFHQ-70k-ADA.
Negative Model Baseline Yu-2019 AuthPrint
ASR Steps LPIPS ASR Steps LPIPS ASR Steps LPIPS
FFHQ-1k (ADA aug.) 100.0 2.17 0.0016 100.0 12.98 0.0076 0.0
FFHQ-30k (ADA aug.) 100.0 1.06 0.0002 100.0 4.52 0.0008 0.0
FFHQ-70k (BCR aug.) 100.0 1.23 0.0001 100.0 3.54 0.0003 0.0
FFHQ-70k (no aug.) 100.0 1.45 0.0001 100.0 3.02 0.0003 0.0
Results show that AuthPrint is resistant to this attack: the attacker fails to generate even a single successful forgery. As shown in Table I, both baseline detectors are defeated by PGD, with a 100% attack success rate—i.e., all non-authentic images are misclassified as authentic. Moreover, the attacks maintain LPIPS consistently below $0.01$ for the baseline detectors, meaning that the attacks are visually imperceptible. While the Yu-2019 detector requires slightly more steps to evade, both are ultimately ineffective. In contrast, AuthPrint achieves a 0% attack success rate. The attacker’s inability to access the secret pixel indices or the reconstructor’s parameters prevents them from training a viable surrogate. This secrecy is vital for the security of AuthPrint: if the attacker had white-box access to the reconstructor, or knowledge of the fingerprint index vector, a PGD attack would likely succeed.
V-F2 Fingerprint Recovery Attacks
Beyond evasion-based attacks, an adversary might also seek to recover the fingerprint indices $s$ in order to forge outputs. We decompose this attack into two steps: first, recovering the fingerprint set $S:=\mathrm{set}(s)$; second, searching for the ordered vector $s$ from $S$. Both the set and the order are crucial—without any one of them, the adversary cannot replicate the behavior of $R$.
Step 1: Recover the set $S$. To recover the set $S$, one can probe the detector through pixel manipulation: starting from an authentic image $x\in\mathbb{R}^{d}$, the attacker manipulates a subset $M\subset[d]$ of $m$ pixels (e.g., setting them to zero) to produce a modified image $x^{\prime}$, and observes whether it passes or is rejected by the verifier. The detector of AuthPrint is designed to compute $D(x;R,s)=\mathrm{MSE}[R(x),x_{s}]$, which remains low upon manipulation if two conditions hold: (1) the reconstructor output $R(x)$ is stable to the manipulation, and (2) the manipulated pixels $M$ do not significantly overlap with the fingerprint set $S$.
1.1: Reconstructor Stability Assumption. Throughout the analysis, we assume that $R(x)$ is stable under small manipulations. To validate this empirically, we generate 100 authentic images using a StyleGAN2 model (FFHQ-70k-ADA), and apply 1,000 random manipulation trials per image. For each trial, we randomly select $m\in\{1,2,4,\ldots,65536\}$ pixel indices to modify, and compute the mean $\ell_{1}$ distance between the reconstructor outputs $R(x)$ and $R(x^{\prime})$. Results in Figure 9 confirm that reconstructor outputs remain stable for small values of $m$.
[…]
1.2: Detector Rejection Criterion. We then formalize the detector’s rejection criterion upon pixel manipulation by analyzing the change in MSE between $x$ and $x^{\prime}$. Let $l=|S|$ be the fingerprint length, and let $\delta^{2}:=\mathbb{E}[(x_{i}-x^{\prime}_{i})^{2}]$ denote the expected pixel perturbation variance over manipulated indices. Under reconstructor stability assumption (i.e., $R(x^{\prime})\approx R(x)=r$), the change in MSE due to pixel manipulation is:
$\Delta\mathrm{MSE}(x^{\prime},x)=\frac{1}{l}\sum_{i\in S}\left[(r_{i}-x^{\prime}_{i})^{2}-(r_{i}-x_{i})^{2}\right].$ (1)
Since only manipulated fingerprint pixels affect the MSE, we can simplify Eq. 1:
$\Delta\mathrm{MSE}(x^{\prime},x)=\frac{1}{l}\sum_{i\in M\cap S}\left[(x_{i}-x^{\prime}_{i})^{2}\right]\approx\frac{|M\cap S|\cdot\delta^{2}}{l}.$ (2)
The detector rejects only if the change in MSE exceeds a fixed threshold $\tau$. Thus, the minimum number of manipulated fingerprint pixels needed for rejection is:
$|M\cap S|\geq k:=\left\lceil\frac{\tau l}{\delta^{2}}\right\rceil.$ (3)
This quantity $k$ defines the rejection threshold—the smallest overlap size between $M$ and $S$ needed to trigger rejection. As a representative example from our evaluation, with $l=32$, $\tau=0.01$, and $\delta^{2}=0.25$, the calculation yields $k=\lceil 1.28\rceil=2$.
The problem of learning a subset of targeted pixels by systematically querying subsets of pixels and receiving one bit of information if the subset contains at least $k$ targeted pixels mirrors the non-adaptive threshold group testing (NATGT) problem [30]: the secret fingerprint set $S$ corresponds to the hidden “defective” set in NATGT, and each attacker query corresponds to NATGT’s test set $M_{i}$. The detector returns a binary signal—accept or reject—based on whether $|M_{i}\cap S|\geq k$, where $k$ is known as the threshold. In NATGT, the goal is to find the minimum number of queries that would allow to identify the defective set, which corresponds to the goal of the attacker in identifying the set $S$, which as few queries to the detector as possible.
1.3: Theoretical Query Complexity. Theoretical analysis of NATGT provides an upper bound on query complexity $T$ to recover $S$:
$T=\mathcal{O}\left(\left(\tfrac{l+1}{k}\right)^{k}\left(\tfrac{l+1}{l-k+1}\right)^{l-k+1}(l+1)\ln\left(\tfrac{d}{l+1}\right)\right),$ (4)
as derived from [31]. For example, with $d=3\cdot 256^{2}$, $l=32$, and $k=2$, this yields $T=\mathcal{O}(10^{5})$ queries. Such query volume is already impractical if the verifier enforces strict budgets for authenticated accounts to the system, and API restrictions like throttling queries per user.
Step 2: Recover the vector $s$. Even if the adversary succeeds in recovering the unordered set $S$, the attack remains ineffective without knowledge of the ordered fingerprint vector $s$. This is because the reconstructor $R$ is trained to predict fingerprint values at positions specified by the exact ordering of $s$; any mismatch prevents training a surrogate $\hat{R}$ that can faithfully mimic $R$ on out-of-distribution inputs. The overall query complexity is therefore the cost of recovering $S$ in Step 1 plus the cost of recovering the ordering in Step 2.
Recovering $s$ from $S$ requires brute-force search over all permutations of $S$, which has factorial complexity $O(l!)$. For example, for a fingerprint length of $l=32$, this corresponds to approximately $O(10^{35})$ permutations. Thus, for sufficiently large fingerprints, the complexity of Step 2 quickly becomes prohibitive and typically dominates the attack cost, far exceeding that of Step 1. As shown in our experiments with StyleGAN2 and SD on $\times256$ RGB images, the optimal fingerprint length is at least $l=16$, already rendering the attack impractical for any realistic adversary. Attacks against models for small images, such as (e.g., $\times8$), where the complexity of both steps may be feasible; however, such resolutions are unrealistic for modern generative models, which continue to strive toward higher image dimensions.
Ensemble Robustness. Robustness against fingerprint recovery can be further improved via an ensemble of detectors. If the verifier accepts an image when any detector in the ensemble accepts it, the adversary must avoid triggering all detectors simultaneously to succeed. This ensemble strategy increases robustness by reducing FNR and increasing the minimum number of overlapping pixels needed for rejection across detectors. Note that this stands in contrast to max-MSE ensembles aimed at minimizing FPR.
VI Discussion and Conclusion
As generative models are increasingly deployed in high-stakes domains such as healthcare and the legal system, ensuring accountability and maintaining the integrity of certified models after deployment becomes critical. Regulatory frameworks, such as the EU AI Act, underscore the importance of independent audits of machine learning systems. Yet, existing approaches often fail to address the risk of a provider substituting a certified model with a cheaper or lower-quality alternative post-certification. This work has introduced AuthPrint, a passive black-box fingerprinting framework designed to verify whether an image originates from a certified generative model, even in settings where the provider may act adversarially.
VI-A Merits of AuthPrint
AuthPrint advances the state of the art along several dimensions. First, it directly addresses the open-world binary verification problem: determining whether an image was generated by a specific model, rather than classifying among a fixed set of known models, as in prior fingerprinting approaches. This distinction makes it suitable for real-world audit scenarios in which the integrity of a single certified model must be preserved. Second, it operates in a fully black-box setting, requiring no access to model internals, specialized hardware, or parameter modifications (as typically required in watermarking), thereby ensuring broad deployability. Third, existing fingerprinting techniques assume honest model providers and focus on preventing fingerprint removal by external attackers. In contrast, AuthPrint addresses a threat model where the model provider itself may attempt to deceive the fingerprinter—a threat model that had not yet been explored in prior work. To defend against a malicious model provider, AuthPrint employs output-dependent fingerprints tied to secret pixel locations known only to the verifier. By leveraging principles from out-of-distribution detection, this design captures subtle distributional shifts caused by model substitution, while the secrecy of fingerprint indices provides robustness against adaptive adversarial strategies, including evasion and reverse-engineering.
We evaluate AuthPrint on two widely used model families—StyleGAN2 and SD—under diverse substitution strategies, including optimization techniques like quantization, pruning, and post-processing. Experiments show that AuthPrint achieves perfect detection performance on GAN models, with FPR@95%TPR reduced to zero. Diffusion models present additional challenges due to their iterative denoising process; however, we demonstrate that practical strategies, such as scaling training data and reconstructor capacity, and employing ensemble detectors, yield high detection performance on SD models, comparable to StyleGAN2.
From a security perspective, AuthPrint exhibits strong resilience against adaptive adversarial attacks. Gradient-based evasion techniques (e.g., PGD) are ineffective because adversaries lack access to the reconstructor and fingerprint indices. For fingerprint recovery, we reduce the task to group testing and show that recovering the complete set of fingerprinted pixels requires an infeasible number of queries; moreover, reconstructing their precise ordering is even more difficult due to the high query complexity of the attack strategy. These results highlight the robustness of AuthPrint when its internal details remain hidden from the provider.
A caveat lies in the secrecy assumptions underlying $R$. While secrecy of the reconstructor is central to AuthPrint, the practical feasibility of upholding this assumption deserves scrutiny. Our prototype uses a 32M-parameter reconstructor for StyleGAN2 and a 674M-parameter reconstructor for SD, but larger reconstructors may be required to verify high-capacity generators such as the latest SD under diverse prompts. Larger models raise operational risks (e.g., expanded hardware attack surfaces or potential leakage of parameters through infrastructure vulnerabilities) as well as scalability challenges, since the auditor must maintain a distinct $R$ for each certified model. These considerations do not undermine the conceptual validity of AuthPrint, but they highlight deployment trade-offs and the need for trusted, well-resourced third-party verifiers capable of safeguarding $R$ throughout the model’s lifetime.
VI-B Limitations of AuthPrint
Nonetheless, certain limitations remain. Detection accuracy may degrade with highly diverse prompts in conditional models such as SD, and quantization is more difficult to detect than other compression techniques such as pruning. These challenges may be mitigated through improved computational strategies during reconstructor training or through pipeline optimizations, pointing to an important avenue for future research. Extensions of AuthPrint could also target other modalities—including audio, video, and text—or explore fingerprinting in alternative domains such as frequency or hybrid spatial–frequency representations. Finally, integrating cryptographic techniques, such as zero-knowledge proofs or trusted execution environments, could enable powerful hybrid verification frameworks.
VI-C Conclusion
In conclusion, AuthPrint provides the first black-box ML-based method for verifying whether generated images originate from a certified model under a threat model that treats the provider as potentially adversarial. It addresses the open-world verification setting, requires no access to model internals, and demonstrates robustness against a range of adaptive attack strategies. Our evaluation across GAN and diffusion models highlights both the effectiveness of the approach and the remaining challenges, such as diverse prompts and quantization. Overall, AuthPrint offers a promising foundation for auditing generative models and motivates further work on extending verification across modalities and strengthening robustness in real-world deployment scenarios.
VII Usage of LLMs
LLMs were used for editorial purposes in this manuscript, and all outputs were inspected by the authors to ensure accuracy and originality.